Rhonda deposits $5,600 in a savings account that pays 2 1/2% interest,compounded semiannually.
A.how much interest does the account earn in the first six months?
B.what is the ending balance after six months ?
C. How much interest does the account earn in the second six months?
D.what is the balance after one year ?
E.how much interest does the account earn the first year?

Answer by rfer(16322)

A)
I=Prt
I=5600*0.025*0.5
I=$70
B)
$5600+705670
C)
A=5600(1+0.025/2)^2
A=5600(1.0125)^2
A=5600*1.02515625
A=5740.88
Interest for year=140.88
140.88-70=$70.88
D)
$5740.88
E)
$140.88